OpenStudy (anonymous):

How many possible positive, negative and complex solutions: x^3−8x^2+15x−56=0

OpenStudy (anonymous):

\[\left\{\{x\to 7\},\left\{x\to \frac{1}{2} \left(1-i \sqrt{31}\right)\right\},\left\{x\to \frac{1}{2} \left(1+i \sqrt{31}\right)\right\}\right\} \]One positive real and two complex solutions.